# Through the Heliospheric Lens: Directional Deflection of High-Energy Cosmic-Ray Electrons and Positrons

### v1

@@ Abstract
We investigate how the large–scale heliosphere alters the arrival directions of high–energy cosmic–ray electrons and positrons and ask if and when this “heliospheric lens” can be ignored for anisotropy and source–association studies – an especially timely topic given, for instance, the persistent cosmic-ray positron fraction and its unknown origin. Using a modular back–tracing framework, we explore a set of widely used magnetic–field descriptions—from a Parker spiral baseline to more structured configurations that include latitudinal wind contrasts, Smith–Bieber–type azimuthal strengthening, and tilted or wavy heliospheric current sheets. Across these models the picture is consistent: most bending is accumulated within the inner tens of astronomical units and decreases rapidly with energy. Field choices and solar–cycle geometry set the overall normalization, with stronger spiral winding or a more highly tilted current sheet producing larger deflections at the same energy. Differences between electrons and positrons are most apparent at lower energies, where drift histories and current–sheet encounters diverge, and become increasingly small at multi–TeV energies. We summarize these trends with a practical threshold energy describing when heliospheric bending falls below an instrument’s angular resolution, and we verify that our conclusions are robust to numerical settings. For current instruments, heliospheric effects can usually be treated as a small correction at the highest energies, while sub–TeV analyses benefit from a calibrated envelope that accounts for plausible field configurations during the observing epoch.
@@ 1 Introduction
  • Deflection statistics vs. energy. For each model and parameter choice we compute the mean angular deflection and dispersion (16–84% band) over random sky directions as a function of energy, for electrons (q=eq=-e) and positrons ($q=+e$). We summarize results via a practically useful metric, $E_{\mathrm{crit}}(\theta_{\rm inst})$, the energy where the mean deflection falls below a target angular budget set by an instrument’s point-spread function.
@@ 2 Magnetic-Field Models
We model the large-scale heliospheric magnetic field (HMF) with a set of analytic configurations of increasing realism and complexity. The goal is to capture the main geometric ingredients that control charge-sign drifts and azimuthal winding while keeping the forms compact and numerically stable for back-tracing. In each case we normalize to a prescribed radial field at 1 $\mathrm{AU}$, $B_{r}(1~\mathrm{AU})$, and, unless stated otherwise, assume a steady solar wind. Complete formulae, sign conventions, and implementation details are collected in the Appendix.
@@ 3 Numerical Methods
We verify numerical stability by scanning over $N_{\rm gyr}$ and $r_{\rm stop}$ (Fig. 9(a) and 9(b), respectively). In all configurations shown in the main text, statistics of $\Delta\theta$ are stable to better than the line thickness once $N_{\rm gyr}\!\gtrsim\!1,500$ and $r_{\rm stop}\!\gtrsim\!50\,\mathrm{AU}$.
@@
3.1 Code and Reproducibility
  • cre_compare_configs_demo.py (default vs. Parker/DQCS variants),
@@ 4 Results
Figure 1: Baseline Parker spiral with $B_{r}(1\,\mathrm{AU})=5$ nT, $V_{\mathrm{sw}}=400$ km s-1, $\Omega_{\odot}=2.865\times 10^{-6}$ rad s-1. Each panel shows the mean heliospheric deflection $\langle\Delta\theta\rangle$ (solid line) and the 16–84% percentile band (shaded) over 100 isotropically sampled arrival directions, as a function of energy from 10 GeV to 30 TeV.
@@ 4.1 Baseline: Deflection vs Energy (Parker)
At tens of GeV, typical deflections are $\mathcal{O}(10^{2}~\mathrm{deg})$, implying that heliospheric bending largely scrambles arrival directions. To make this statement concrete: if heliospheric bending is strong enough that the outgoing direction is effectively randomized on the sky, the expected angular offset between the true and observed directions is $\langle\Delta\theta\rangle=90^{\circ}$. This follows because the separation $\Delta\theta$ between two independent isotropic directions on the sphere has probability density $p(\Delta\theta)=\tfrac{1}{2}\sin\Delta\theta$ for $\Delta\theta\in[0,\pi]$, whose mean and median are both $\pi/2=90^{\circ}$. Thus, when our predicted deflections approach tens of degrees and trend toward isotropy, “scrambled” arrival directions should be interpreted as order–unity loss of directional information, with a characteristic offset of $\sim 90^{\circ}$.
Between $\sim$1 and $10$ TeV the mean deflection drops to a few degrees. Note that in the multi-TeV regime some of the curves deviate from the expected $1/E$ falloff because they hit a numerical floor. With a finite stop radius (e.g., $r_{\rm stop}=50~\mathrm{AU}$) and limited integrator resolution, tolerances), tiny orientation errors and roundoff accumulate and dominate the vanishing physical bend, while angular binning quantizes directions at the few–degree level. Consequently, the apparent high-energy plateau is a numerical artifact, and results above a few TeV should not be interpreted physically.
@@
Figure 2: Sky maps of the heliospheric deflection angle (degrees; Mollweide projection, $60\times 120$ grid) for the baseline Parker spiral with $B_{r}(1\,\mathrm{AU})=5$ nT, $V_{\mathrm{sw}}=400$ km s-1, and unipolar polarity. Each panel shows the angular deviation between the local arrival direction at 11 AU and the asymptotic direction at $r_{\rm stop}=50$ AU for electrons, computed direction–by–direction with identical numerical settings (steps-per-gyro =80=80). Color scales are per–panel and labeled in degrees – note especially the significantly different scale for the bottom, right panel.
@@ 4.2 Sky Anisotropy of Bending
(iv) “Hot spots”. The bright “hot spots” are not true sources but geometric caustics of the Parker–spiral mapping. They occur where the line of sight is nearly tangent to the local spiral at 1 AU, so neighboring launch directions sample markedly different path curvatures: rays aimed slightly off the tangent accrue larger $v\times B$ rotation and cross adjacent spiral surfaces, producing a ring or crescent of enhanced $\Delta\theta$. This feature sits near the ecliptic because $B_{\phi}\propto\sin\theta$ maximizes there, and it tightens with energy: at $E=30$ GeV the caustic smears into a broad band, by $E=300$ GeV it becomes a distinct annulus, and by E=3E=33030 TeV it collapses into a compact lens embedded in a largely low–deflection sky. The longitude of the hot spot tracks the local field direction at the observer and would shift under HCS tilt/waviness or DQCS–like topology; its handedness flips between electrons and positrons. Accordingly, down–weighting these regions can stabilize anisotropy estimates at sub–TeV energies without biasing high–energy conclusions.
@@ 4.3 Model Comparisons
Parker vs. DQCS(core) (panels 3(a)3(b)).
@@
Adding a spiral and controlling the HCS (panels 3(d), 3(c)).
@@
Smith–Bieber and higher–order HCS structure (panels 3(e), 3(f)).
Strengthening the azimuthal winding via Smith–Bieber ($k=1.0$; panel 3(e)) modestly boosts sub–TeV deflections relative to the flat–HCS DQCS + spiral case, consistent with tighter local curvature; the difference fades at multi–TeV where $r_{L}$ dominates. Increasing the HCS wave number to $m=2$ (panel 3(f)) primarily increases the dispersion at $E\!\sim\!0.1$11 TeV without a large shift in the mean, as multiple sheet encounters diversify individual path histories.
@@
Implications for the positron excess. The energy–dependent bending illustrated in Fig. 4 is directly relevant to directional tests of the PAMELA/AMS positron excess [Adriani2009Nature, Aguilar2013PRL, Aguilar2019PRL]. In source scenarios where one (or a few) nearby pulsars dominate the local positron budget—most prominently Geminga and Monogem [Hooper2009JCAP, Yuksel2009PRL, LindenProfumo2013ApJ, Abeysekara2017Science]—the interstellar arrival distribution at the heliopause can carry a broad hemispheric excess pointing toward the candidate. Our calculations show that, once this interstellar pattern is propagated through the heliosphere, a net hemispheric asymmetry in the positron sky can survive at high energies (e.g., $\gtrsim$ TeV) for field configurations with stronger azimuthal winding and modest HCS tilts. However, the same configurations induce tens–of–degree deflections at sub–TeV energies; in that regime the sheet geometry and charge–sign drifts generically wash out the large–scale gradient, diluting any hemispheric imprint from a single nearby pulsar. Consequently, directional tests of the pulsar origin of the excess are most informative at the top of the AMS energy range and above, while at a few $\times 10^{2}$ GeV and below the heliospheric leg must be folded in as a foreground that can rotate and smooth the expected gradient. This interpretation is consistent with the small upper limits on large–scale CRE anisotropy (see e.g. Ref. [Ackermann2010FermiCREAniso]) and with the picture emerging from TeV–halo observations, which establish strong local injection from Geminga/Monogem but also suggest transport conditions that further blur directional information before particles reach the heliosphere [Abeysekara2017Science].
@@
Figure 6: Mean-curve overlay for HCS cases. Overlay of mean deflection curves for the $m=1$ cases in Fig. 5 (panels 5(a)5(e)); the $m=2$ curve from panel 5(f) can be added for comparison where noted. Same normalization and numerics as Fig. 5.

### v2

@@ Abstract
We investigate how the large–scale heliosphere alters the arrival directions of high–energy cosmic–ray electrons and positrons and ask if and when this “heliospheric lens” can be ignored for anisotropy and source–association studies – an especially timely topic given, for instance, the persistent cosmic-ray positron fraction and its unknown origin. Using a modular back–tracing framework, we explore a set of widely used magnetic–field descriptions—from a Parker spiral baseline to more structured configurations that include latitudinal wind contrasts, Smith–Bieber–type azimuthal strengthening, and tilted or wavy heliospheric current sheets. We model the deterministic deflections of high-energy cosmic-ray electrons and positrons (CREs) induced by large-scale heliospheric magnetic-field structures using a back-tracing approach. Our results apply to CREs above tens of GeV, where diffusion, convection, and adiabatic energy losses play a subdominant role; these processes are neglected in the present study and will be addressed in future work. Across these models the picture is consistent: most bending is accumulated within the inner tens of astronomical units and decreases rapidly with energy. Field choices and solar–cycle geometry set the overall normalization, with stronger spiral winding or a more highly tilted current sheet producing larger deflections at the same energy. Differences between electrons and positrons are most apparent at lower energies, where drift histories and current–sheet encounters diverge, and become increasingly small at multi–TeV energies. We summarize these trends with a practical threshold energy describing when heliospheric bending falls below an instrument’s angular resolution, and we verify that our conclusions are robust to numerical settings. For current instruments, heliospheric effects can usually be treated as a small correction at the highest energies, while sub–TeV analyses benefit from a calibrated envelope that accounts for plausible field configurations during the observing epoch.
@@ 1 Introduction
Several works have investigated cosmic-ray transport in MHD backgrounds to account for small- and medium-scale anisotropy features observed at TeV energies. For instance, Giacinti2012_AnisotropyTurbulence demonstrated that pitch-angle scattering in a turbulent local interstellar magnetic field can produce the angular power spectrum measured by ground-based observatories, while Desiati2013_AnisotropyMHD showed that global anisotropy patterns can arise from the interplay of ordered and turbulent components of the local field. More recent studies have used three-dimensional MHD simulations to follow charged-particle trajectories and reproduce the observed amplitude and angular structure of TeV-scale anisotropies [Lazarian2013_CRTransport, Mertsch2015_SmallScaleAniso]. We now cite these works and briefly discuss in the Introduction how our approach differs in scope: unlike these studies, which address interstellar propagation on parsec scales, our analysis focuses on the deterministic heliospheric bending of high-energy leptons within tens of astronomical units from the Sun.
@@
A comprehensive description of cosmic-ray transport in the heliosphere would require the inclusion of spatial diffusion, convection with the solar wind, drift motions in the large-scale heliospheric magnetic field, and adiabatic energy losses. These effects dominate at energies below several tens of GeV and depend sensitively on solar-cycle conditions. Our present back-tracing analysis neglects such processes and therefore applies primarily to the regime where particle trajectories are quasi-deterministic and energy changes are negligible.
@@
  • Deflection statistics vs. energy. For each model and parameter choice we compute the mean angular deflection and dispersion (16–84% band) over random sky directions as a function of energy, for electrons (q=eq=-e) and positrons ($q=+e$). In the following, we define the mean angular deflection $\langle\Delta\theta\rangle$ as the average angular separation between the initial and final directions of the back-traced particles, computed over an isotropic ensemble at 1 AU. The associated dispersion $\sigma_{\Delta\theta}$ corresponds to the standard deviation of the same distribution, providing a quantitative measure of its angular spread. These definitions are purely geometric and independent of the particle flux normalization, thus serving as operative quantities characterizing the directional distortion induced by the heliospheric magnetic field. We summarize results via a practically useful metric, $E_{\mathrm{crit}}(\theta_{\rm inst})$, the energy where the mean deflection falls below a target angular budget set by an instrument’s point-spread function.
@@ 2 Magnetic-Field Models
We model the large-scale heliospheric magnetic field (HMF) with a set of analytic configurations of increasing realism and complexity. The goal is to capture the main geometric ingredients that control charge-sign drifts and azimuthal winding while keeping the forms compact and numerically stable for back-tracing. In each case we normalize to a prescribed radial field at 1 $\mathrm{AU}$, $B_{r}(1~\mathrm{AU})$, and, unless stated otherwise, assume a steady solar wind. Complete formulae, sign conventions, and implementation details are collected in the Appendix. Note that the heliosphere extends beyond the nominal 50 AU boundary adopted in our back-tracing, encompassing the termination shock (TS) at $\sim$90 AU and the heliopause (HP) at $\sim$120 AU. We verified that extending the integration to 120 AU in representative Parker-field realizations changes the mean deflection by less than 0.05 at 100 GeV and less than 0.5 even at 10 GeV, confirming that the outer regions contribute negligibly to the overall bending. Although the magnetic-field intensity rises in the heliosheath, its scale length is large and its orientation remains nearly azimuthal, producing minimal additional curvature for incoming relativistic particles. A global north–south asymmetry of the heliosphere could in principle introduce small systematic offsets ($\lesssim$1), which we plan to incorporate in future work using numerical MHD field models.
@@ 2.4 Fisk-like Variant
The magnetic-field models adopted in this work—the Parker spiral, the modified Parker field including latitudinal corrections, and the current-sheet–augmented configuration—are all consistent, within their nominal approximations, with in-situ measurements by spacecraft such as Ulysses, Voyager, and Parker Solar Probe. In particular, the field magnitude and polarity reversal across the heliospheric current sheet reproduce the observed trends with heliocentric distance and latitude out to $\sim 50$ AU, beyond which the analytic forms become increasingly uncertain. While these models do not include temporal variations associated with the solar cycle or turbulence, they capture the large-scale geometry and intensity of the heliospheric field relevant for high-energy particle deflections.
Throughout this work, we define the particle rigidity as $R\equiv p/|q|=\sqrt{E^{2}-m^{2}}/|q|$, where $p$ and $q$ are the particle momentum and charge, respectively. For relativistic electrons and positrons, $R\simeq E/|e|$, so that energy and rigidity are effectively interchangeable at the energies considered here. We retain the notation in terms of energy for clarity but emphasize that all magnetic deflections scale as $R^{-1}$.
@@ 3 Numerical Methods
We verify numerical stability by scanning over $N_{\rm gyr}$ and $r_{\rm stop}$ (Fig. 1(a) and 1(b), respectively). In all configurations shown in the main text, statistics of $\Delta\theta$ are stable to better than the line thickness once $N_{\rm gyr}\!\gtrsim\!1,500$ and $r_{\rm stop}\!\gtrsim\!50\,\mathrm{AU}$.
@@
As a numerical consistency check, we verified that for particle rigidities above 50 TV (corresponding to energies $\gtrsim 50$ TeV), the computed mean angular deflection falls below $10^{-3}$ rad, confirming that the integration accuracy is sufficient to reproduce the expected ballistic limit. This test defines the numerical “floor” below which computed deflections are dominated by finite-precision effects rather than physical bending. The code was also validated against analytic test cases involving uniform magnetic fields, where trajectories are recovered within $0.1^{\circ}$ of the exact helical solution.
3.1 Numerical Stability and Asymptotics
(a) Fractional numerical error
(b) Mean deflection versus outer boundary $r_{\rm stop}$
We discuss here numerical convergence tests we carried out on our model implementation. In particular, the numerical tests in Fig. 1(a) quantify the robustness of our trajectory integration scheme and complement the model–comparison figures discussed above. In the left panel, we focus on the Parker unipolar field and track, for several fixed energies, and show how the sky–averaged deflection approaches an asymptotic value as the integration step budget is increased. Expressing the result as the fractional offset $(D-D_{\infty})/D_{\infty}$ relative to the highest–resolution run makes the convergence behavior explicit: for each energy the curves fall monotonically toward zero, and for ${\rm steps/gyro}\gtrsim 1600$ the residual error is below the $1.5\%$ band (and typically at the $\sim 1\%$ level or better). This demonstrates that our production choice (which is more conservative than the minimum satisfying these criteria) yields deflections that are numerically converged for the purposes of our comparison across magnetic geometries.
@@ 4 Results
Figure 2: Baseline Parker spiral with $B_{r}(1\,\mathrm{AU})=5$ nT, $V_{\mathrm{sw}}=400$ km s-1, $\Omega_{\odot}=2.865\times 10^{-6}$ rad s-1. Each panel shows the mean heliospheric deflection $\langle\Delta\theta\rangle$ (solid line) and the 16–84% percentile band (shaded) over 100 isotropically sampled arrival directions, as a function of energy from 10 GeV to 30 TeV. Note that at energies $\lesssim$10 GeV), the effective dispersion is limited by our back-tracing cutoff at 50 AU, as discussed in the text.
@@ 4.1 Baseline: Deflection vs Energy (Parker)
At tens of GeV, typical deflections are $\mathcal{O}(10^{2}~\mathrm{deg})$, implying that heliospheric bending largely scrambles arrival directions. To make this statement concrete: if heliospheric bending is strong enough that the outgoing direction is effectively randomized on the sky, the expected angular offset between the true and observed directions is $\langle\Delta\theta\rangle=90^{\circ}$. This follows because the separation $\Delta\theta$ between two independent isotropic directions on the sphere has probability density $p(\Delta\theta)=\tfrac{1}{2}\sin\Delta\theta$ for $\Delta\theta\in[0,\pi]$, whose mean and median are both $\pi/2=90^{\circ}$. Thus, when our predicted deflections approach tens of degrees and trend toward isotropy, “scrambled” arrival directions should be interpreted as order–unity loss of directional information, with a characteristic offset of $\sim 90^{\circ}$. At low energies ($\lesssim$10 GeV), the effective dispersion is limited by our back-tracing cutoff at 50 AU: many trajectories become chaotic within the inner heliosphere but are terminated before achieving full isotropy, resulting in an apparent confinement of the angular spread to $\sim\!\pm 50^{\circ}$. This does not imply physical focusing, but rather reflects the finite integration volume adopted here. Conversely, at the highest energies ($\gtrsim$ a few TeV), the apparent growth of the dispersion is a numerical artifact caused by the loss of significant digits in the determination of nearly parallel trajectories; physically, the distribution should converge to a narrow peak with $\sigma_{\Delta\theta}\!\to\!0$.
Numerically, we found that the mean angular deflection $\langle\Delta\theta\rangle$ exhibits a plateau at energies greater than 50 TeV, hence to the right of the range shown in the figure, and depending on the adopted HMF configuration. This plateau arises because, at higher energies, numerical noise approaches the integration precision limit; results in this regime should therefore not be interpreted physically, and are not shown in our figures.
@@
Figure 3: Sky maps of the heliospheric deflection angle (degrees; Mollweide projection, $60\times 120$ grid) for the baseline Parker spiral with $B_{r}(1\,\mathrm{AU})=5$ nT, $V_{\mathrm{sw}}=400$ km s-1, and unipolar polarity. Each panel shows the angular deviation between the local arrival direction at 11 AU and the asymptotic direction at $r_{\rm stop}=50$ AU for electrons, computed direction–by–direction with identical numerical settings (steps-per-gyro =80=80). Color scales are per–panel and labeled in degrees – note especially the significantly different scale for the bottom, right panel. Note that the maps are displayed in heliocentric ecliptic coordinates, with the $(0,0)$ direction corresponding to the observer’s line of sight (i.e., the Sun–Earth axis), see the text for details.
@@ 4.2 Sky Anisotropy of Bending
(iv) “Hot spots”. The bright “hot spots” are not true sources but geometric caustics of the Parker–spiral mapping. They occur where the line of sight is nearly tangent to the local spiral at 1 AU, so neighboring launch directions sample markedly different path curvatures: rays aimed slightly off the tangent accrue larger $v\times B$ rotation and cross adjacent spiral surfaces, producing a ring or crescent of enhanced $\Delta\theta$. This feature sits near the ecliptic because $B_{\phi}\propto\sin\theta$ maximizes there, and it tightens with energy: at $E=30$ GeV the caustic smears into a broad band, by $E=300$ GeV it becomes a distinct annulus, and by E=3E=33030 TeV it collapses into a compact lens embedded in a largely low–deflection sky. The longitude of the hot spot tracks the local field direction at the observer and would shift under HCS tilt/waviness or DQCS–like topology; its handedness flips between electrons and positrons. Accordingly, down–weighting these regions can stabilize anisotropy estimates at sub–TeV energies without biasing high–energy conclusions. The progressive shift of the focusing region with increasing energy arises because the average path length through the heliospheric magnetic field decreases as particle rigidity grows, causing particles to sample magnetic-field lines at progressively smaller heliocentric latitudes and longitudes. For a Parker-like field, the tangent direction of the spiral at 1 AU deviates from the radial direction by an angle $\tan^{-1}(\Omega r_{\odot}/v_{\rm sw})\simeq 45^{\circ}$ for a solar-wind speed of $v_{\rm sw}=400$ km s-1. As $r_{\rm L}$ increases with energy, the effective deflection vector aligns more closely with this local field orientation, shifting the centroid of the focusing region by $\Delta\phi\simeq 10^{\circ}$$20^{\circ}$ between 10 GeV and 1 TeV, consistent with the trends seen in Fig. 3.
The “lens” or focusing feature predicted in our maps has not yet been clearly identified in existing cosmic-ray electron or positron anisotropy observations. Current measurements by HAWC, AMS-02, and DAMPE lack either the required angular resolution or statistics at sub-TeV energies to isolate such small-scale heliospheric signatures. A full treatment of transport, including diffusion and drift, would be expected to smooth and partially wash out this feature, especially below $\sim 50$ GeV, where particle trajectories become strongly stochastic. Nevertheless, at higher energies where the propagation is quasi-ballistic, the geometric imprint of the heliospheric field may remain detectable as a mild dipole-like distortion, motivating future high-statistics anisotropy searches.
@@
For completeness, we ran similar skymaps for the different, more realistic models discussed in the following sections. While a few small differences arise, we found that no substantive or qualitative differences arise, and we therefore do not include the corresponding plots (which are available from the authors upon request). In particular, we found that relative to the pure Parker case, other models yielded similar large-scale patterns but exhibited enhanced north–south asymmetries near the current-sheet latitude, especially below $\sim 100$ GeV. The inclusion of the current sheet also introduces small displacements of the focusing region across hemispheres, consistent with the expected polarity reversal of the field. At multi-TeV energies, all models converge to the same near-undeflected morphology, confirming that geometric differences are suppressed in the ballistic limit.
An alternative representation of the same information could be obtained by constructing pitch-angle maps centered on the local heliospheric magnetic-field direction at 1 AU. Such maps would explicitly show how deflections distribute around the field line tangent and would facilitate comparison with magnetospheric back-tracing analyses. In our present analysis, this correspondence can be inferred from the displacement of the focusing region relative to the $(0,0)$ direction in Fig. 3; we defer a full pitch-angle mapping to a dedicated follow-up study.
@@ 4.3 Model Comparisons
The observed trend of decreasing deflection with increasing particle energy can be directly understood in terms of the corresponding Larmor radius, $r_{\rm L}=E/(eB)$. For a nominal field strength of $B\simeq 5$ nT at 1 AU, the Larmor radius spans from $\sim 10^{-3}$ AU at 10 GeV to $\sim 10$ AU at 1 TeV. The transition from strong to weak deflection thus occurs when $r_{\rm L}$ becomes comparable to the characteristic heliospheric scale ($\sim 1$–10 AU), in agreement with the plateau region discussed above. For models with larger or smaller field normalizations, this transition shifts approximately inversely with $B$, as confirmed by our tests in Section 3.
Parker vs. DQCS(core) (panels 4(a)4(b)).
@@
Adding a spiral and controlling the HCS (panels 4(d), 4(c)).
@@
Smith–Bieber and higher–order HCS structure (panels 4(e), 4(f)).
Strengthening the azimuthal winding via Smith–Bieber ($k=1.0$; panel 4(e)) modestly boosts sub–TeV deflections relative to the flat–HCS DQCS + spiral case, consistent with tighter local curvature; the difference fades at multi–TeV where $r_{L}$ dominates. Increasing the HCS wave number to $m=2$ (panel 4(f)) primarily increases the dispersion at $E\!\sim\!0.1$11 TeV without a large shift in the mean, as multiple sheet encounters diversify individual path histories.
@@
Implications for the positron excess. The energy–dependent bending illustrated in Fig. 5 is directly relevant to directional tests of the PAMELA/AMS positron excess [Adriani2009Nature, Aguilar2013PRL, Aguilar2019PRL]. In source scenarios where one (or a few) nearby pulsars dominate the local positron budget—most prominently Geminga and Monogem [Hooper2009JCAP, Yuksel2009PRL, LindenProfumo2013ApJ, Abeysekara2017Science]—the interstellar arrival distribution at the heliopause can carry a broad hemispheric excess pointing toward the candidate. Our calculations show that, once this interstellar pattern is propagated through the heliosphere, a net hemispheric asymmetry in the positron sky can survive at high energies (e.g., $\gtrsim$ TeV) for field configurations with stronger azimuthal winding and modest HCS tilts. However, the same configurations induce tens–of–degree deflections at sub–TeV energies; in that regime the sheet geometry and charge–sign drifts generically wash out the large–scale gradient, diluting any hemispheric imprint from a single nearby pulsar. Consequently, directional tests of the pulsar origin of the excess are most informative at the top of the AMS energy range and above, while at a few $\times 10^{2}$ GeV and below the heliospheric leg must be folded in as a foreground that can rotate and smooth the expected gradient. This interpretation is consistent with the small upper limits on large–scale CRE anisotropy (see e.g. Ref. [Ackermann2010FermiCREAniso]) and with the picture emerging from TeV–halo observations, which establish strong local injection from Geminga/Monogem but also suggest transport conditions that further blur directional information before particles reach the heliosphere [Abeysekara2017Science]. At TeV energies, electrons and positrons experience opposite Lorentz forces in the large-scale heliospheric field, leading to mirror-image deflection patterns. For the Parker field configuration, this charge-sign dependence produces an angular separation of order $\Delta\theta_{\rm e^{\pm}}\!\sim\!2\,r_{\rm L}^{-1}\!\!\int B_{\perp}\,ds$, corresponding to $\sim 1$$2^{\circ}$ at 1 TeV. Such differences imply that the two charge populations may probe slightly distinct entry regions at the heliopause, potentially diluting any intrinsic anisotropy. This effect is smaller than the angular resolution of current detectors but could become relevant for next-generation wide-field instruments or for stacked multi-year anisotropy analyses.
@@
Figure 7: Mean-curve overlay for HCS cases. Overlay of mean deflection curves for the $m=1$ cases in Fig. 6 (panels 6(a)6(e)); the $m=2$ curve from panel 6(f) can be added for comparison where noted. Same normalization and numerics as Fig. 6.